# SurgiATM: A Physics-Guided Plug-and-Play Model for Deep Learning-Based Smoke Removal in Laparoscopic Surgery

### v1

@@ 3 Methods
The proposed method is motivated by an interesting observation in laparoscopic surgical video smoke removal: predictions from most desmoking methods yield a Laplacian-like error distribution (see Fig. 1). Building upon this finding, we introduce a new restoration formula that minimizes prediction errors and enhances model robustness by blending physics-based and deep learning-based approaches from the perspective of probability distributions.
Figure 1: The left images are from a natural dehazing dataset, O-HAZE [Ancuti_2018_O_HAZE]; the middle group is from a real-world surgical desmoking benchmark VASST-desmoke [Xia_2025_In_Vivo]; and the line chart (right) shows the difference in error distributions between surgical desmoking and natural dehazing, computed from the entire dataset. The $\text{1}^{\text{st}}$ row displays the hazy or smoky images; the $\text{2}^{\text{nd}}$ row shows the smoke or haze mask estimated from the ground truth; and the $\text{3}^{\text{rd}}$ row presents the error magnitude map of the DCP prediction.
@@ 3.1.1 Physics-Based Modeling
The atmospheric scattering model [Narasimhan_2002_Vision_and, Narasimhan_2003_Contrast_restoration, Fattal_2008_Single_image, Tan_2008_Visibility_in] is typically formulated as:
@@
where $c$ denotes a color channel; $x$ represents the spatial location of a pixel; $I\in\mathbb{R}^{H\times W\times C}$ is the observed image, where $C$, $W$, and $H$ are the number of color channels, width and height, respectively; $J$ indicates the corresponding haze-free image; $A\in\mathbb{R}^{1\times 1\times C}$ denotes the global atmospheric light; and $t\in\mathbb{R}^{H\times W\times 1}$ is the medium transmission, which can alternatively be expressed as t=1St=1-S with smoke density $S$. In particular, $t$ and $J$ can be further formulated as:
@@
where $\beta(\lambda)\in\mathbb{R}$ is the scattering coefficient for a specific wavelength $\lambda$ of light; $d\in\mathbb{R}^{H\times W\times 1}$ represents the scene depth; and $\rho\in\mathbb{R}^{H\times W\times C}$ is the normalized radiance [Narasimhan_2003_Contrast_restoration] or reflectance [Tan_2008_Visibility_in] of a scene point.
[Kaiming_2011_Single_Image] proposed the Dark Channel Prior (DCP), which estimates the transmission map $t$ using the dark channel:
$\begin{split}t(x)&=1-D(x)\\ &=1-\underset{c\in C}{\mathrm{min}}\left(\underset{u\in\Omega(x;z)}{\mathrm{min}}\left(\frac{I(u,c)}{A(c)}\right)\right),\end{split}$ (4)
@@ 3.1.2 Deep Learning-Based Modeling
Clearly, in pixels where the smoke density is high and white or gray surgical instruments are present (i.e., $D\rightarrow 1$), the conventional physics-based model (5) may fail. Moreover, radial light attenuation can violate the assumption that $A\in\mathbb{R}^{1\times 1\times C}$; under such conditions, $A$ is more appropriately represented as $A\in\mathbb{R}^{H\times W\times C}$. Adhering to the original assumption may introduce more errors, whereas adopting the latter formulation significantly increases the degrees of freedom of the solution space. Therefore, Deep Neural Network (DNN), a powerful technique for non-linear mapping, is widely employed for smoke removal. Its formulation and training objective are typically expressed as:
@@
where $\Phi$ indicates a non-linear mapping parameterized by trainable weights $\theta$; $\theta^{*}$ represents the optimized weights obtained through training on the dataset $\mathcal{X}$; $\tilde{Z}$ and $\mathcal{L}$ denote priors and loss function, respectively, as determined by the specific method design. Nevertheless, its performance and generalizability may be limited when deployed in unseen surgical scenarios.
@@ 3.2 Mixture-of-Experts and Optimization
Inspired by the Mixture-of-Experts (MoE), an effective mechanism for complementary advantages [Jacobs_1991_Adaptive_Mixtures, Ding_2025_DenseFormer_MoE], we begin with an ideal gating model:
$\begin{split}J_{\mathrm{MoE}}(x,c)&=\tilde{W}(x)\cdot J_{\mathrm{DCP}}(x,c)\\ &\quad+(1-\tilde{W}(x))\cdot J_{\mathrm{DNN}}(x,c),\end{split}$ (8)
@@
$\varepsilon_{\mathrm{MoE}}(x,c)$ $=W(x)\cdot\varepsilon_{\mathrm{DCP}}(x,c)$
$\quad+(1-W(x))\cdot\varepsilon_{\mathrm{DNN}}(x,c).$ (12)
@@
$\begin{split}\min_{W(x)}\;\;&\mathbb{E}[\varepsilon_{\mathrm{MoE}}^{2}],\quad s.t.\;\;0\leq W(x)\leq 1.\end{split}$ (13)
Table 1: Comparison of Prediction Error Distributions with Gaussian and Laplacian Distributions via Jensen-Shannon (JS) Divergence in VASST-desmoke with real-world smoke and Cholec80 with synthetic smoke
@@
DCP [Kaiming_2011_Single_Image] 0.044989 0.013664 0.002782 0.010886
AODNet [Li_2017_AOD_Net] 0.077348 0.056151 0.006405 0.012263
CGAN-DC [Salazar_Colores_2020_Desmoking_Laparoscopy] 0.033480 0.033894 0.007939 0.005278
De-smokeGCN [Chen_2020_De_smokeGCN] 0.027138 0.001988 0.007484 0.005802
GCANet [Chen_2019_Gated_Context] 0.009448 0.003621 0.003164 0.008474
LGUTransformer [Wang_2024_Desmoking_of] 0.010483 0.003129 0.016931 0.003623
MARS-GAN [Hong_2023_MARS_GAN] 0.024441 0.064570 0.009878 0.005369
MSBDN [Dong_2020_Multi_Scale] 0.020062 0.007690 0.005995 0.005536
RSTN [Wang_2023_Surgical_smoke] 0.026959 0.006900 0.012523 0.004193
SSIM-PAN [Sidorov_2020_Generative_Smoke] 0.018315 0.011999 0.005900 0.006517
SVPNet [Wang_2024_Smoke_veil] 0.022240 0.000675 0.007876 0.006492
@@
Figure 2: The Laplacian parameters (i.e., $\mu$ and $b$) estimated from the specific dataset, and indicator $W^{*}$ calculated with (19), are presented across different methods with the Confidence Interval (CI). The upper and lower rows correspond to Cholec80 and VASTT-desmoke, respectively.
@@
where $\mathcal{M}$ represents a method, either DCP or DNN. Collecting (12), and (14), we have:
$\begin{split}\mathrm{Var}[\varepsilon_{\mathrm{MoE}}]&=2\cdot\big(W^{2}(x)\cdot b^{2}_{\mathrm{DCP}}(D(x))\\ &\quad+(1-W(x))^{2}\cdot b^{2}_{\mathrm{DNN}}(D(x))\big),\end{split}$ (15)
$\begin{split}\mathbb{E}[\varepsilon_{\mathrm{MoE}}]&=W(x)\cdot\mu_{\mathrm{DCP}}(D(x))\\ &\quad+(1-W(x))\cdot\mu_{\mathrm{DNN}}(D(x)).\end{split}$ (16)
Then, the original optimization problem (13) is rewritten as:
@@
We solve (17) and obtain the unconstrained solution:
@@
where $W^{*}(x;D)$ is the optimized $W(x)$ conditioned on the dark channel $D$ without constraints. For clarity, $b_{\mathrm{DCP}}(D(x))$ and $\mu_{\mathrm{DNN}}(D(x))$ are written as $b_{1,x}$ and $\mu_{2,x}$, respectively. Subsequently, $W^{*}(x)$ is clipped to the interval $[0,1]$ to satisfy the constraint (18).
@@ 3.3 A New Restoration Formula
Following (19), Fig. 2 presents the trends of $W^{*}$ given $D$, based on statistics across all methods in two datasets: Cholec80 [Twinanda_2017_EndoNet_A] with synthetic smoke, and the benchmark VASST-desmoke [Xia_2024_A_New, Xia_2025_In_Vivo] containing real-world smoke. Deep learning models display varied trends, while a consistent negative correlation between $W^{*}$ and $D$ is observed. Therefore, in this work, $W$ is estimated by:
@@
  2. 2.
    Equation (20) is intuitive: a high intensity of dark channel (i.e., $D(x)\rightarrow 1$) statistically indicates dense smoke [Kaiming_2011_Single_Image, Xia_2025_In_Vivo], which challenges DCP due to information loss and division by zero or very small values, thereby making DNN the dominant model (i.e., $W(x)\rightarrow 0$);
@@
  4. 4.
    The denominator 1D1-D and the unknown parameter $A$ in (5) are eliminated in the subsequent derivation, stabilizing the computation, simplifying the modeling, and precluding hard value clipping.
By combining (5), (8), and (20), we obtain:
$\begin{split}J_{\mathrm{MoE}}(x,c)&=I(x,c)-A(c)\cdot D(x)\\ &\quad+D(x)\cdot J_{\mathrm{DNN}}(x,c),\end{split}$ (21)
where $A$ remains an unknown parameter that requires an additional model for its estimation. To address this issue, we first use (3) to further decompose $J_{\mathrm{DNN}}$:
$\begin{split}J_{\mathrm{MoE}}(x,c)&=I(x,c)-A(c)\cdot D(x)\\ &\quad+D(x)\cdot A(c)\cdot\rho_{\mathrm{DNN}}(x,c),\end{split}$ (22)
where $\rho_{\mathrm{DNN}}(x,c)=J_{\mathrm{DNN}}(x,c)/A(c)$ represents the reflectance calculated from the normalized DNN prediction. Equation (22) can be reformulated as:
$\begin{split}J_{\mathrm{MoE}}(x,c)&=I(x,c)-D(x)\cdot A(c)\cdot(1-\rho_{\mathrm{DNN}}(x,c))\\ &=I(x,c)-\mathcal{D}(x)\cdot(1-\rho_{\mathrm{DNN}}(x,c)),\end{split}$ (23)
where $\mathcal{D}$ denotes the denormalized dark channel $D$. Recalling (4), $\mathcal{D}$ is defined as:
@@
where $A$ is eliminated, thereby avoiding the challenging estimation of global light illuminance in laparoscopic surgical scenes. In this case, the original DNN modeling in (6) and (7) can be improved as:
@@
where $\sigma$ is a differentiable function (typically a sigmoid) that maps the DNN output to the interval $[0,1]$. It can be ignored if the output is already constrained within this range, depending on the specific model design. The proposed formulation introduces only minimal modifications to the original DNN model $\Phi_{\theta}$, without adding any additional trainable parameters. Notably, SurgiATM leverages the denormalized dark channel as a key element, in line with recent methods [Chen_2019_Gated_Context, Hong_2023_MARS_GAN, Chen_2020_De_smokeGCN], suggesting that the dark channel prior plays a useful role in surgical desmoking [Xia_2025_In_Vivo].
@@ 3.4 Gradient Computation and Refinement
In this section, we compute the gradient with respect to the model prediction $\varrho_{\mathrm{DNN}}$ for two commonly used loss functions: Mean Absolute Error (MAE or L1) loss and Mean Squared Error (MSE or L2) loss. Their gradients are denoted as $\nabla_{\varrho}\mathcal{L}_{1}$ and $\nabla_{\varrho}\mathcal{L}_{2}$, respectively:
$\nabla_{\varrho}\mathcal{L}_{1}=\begin{cases}-\mathcal{D},&J>1-\mathcal{D}\cdot(1-\rho);\\ \mathcal{D},&J<1-\mathcal{D}\cdot(1-\rho);\\ 0,&J=1-\mathcal{D}\cdot(1-\rho),\end{cases}$ (27)
$\nabla_{\varrho}\mathcal{L}_{2}=-\mathcal{D}\cdot(J-(1-\mathcal{D}\cdot(1-\rho))),$ (28)
@@
where $\eta$ is a smoothing factor; $\eta=0$ corresponds to no refinement; and $\eta\rightarrow+\infty$ implies $\hat{\mathcal{D}}\rightarrow 1$, causing (26) to degenerate into conventional residual prediction. Ablation results for $\eta$ and $z$ are reported in Section 4.5.

### v2

@@ 3 Methods
The proposed method is motivated by an intriguing observation in laparoscopic surgical video smoke removal: predictions from most desmoking methods yield a Laplacian-like error distribution (see Fig. 2). Building upon this insight, we develop a new restoration formula that minimizes prediction errors and enhances model robustness by revisiting classic physical model, analyzing tge predictions of deep learning-based approaches, leveraging a MoE model, and optimizing from the perspective of probability distributions (see Fig. 1). SurgiATM can be seamlessly embedded at the output stage of deep learning methods, where it predicts a normalized radiance used for reconstruction. Integration with general DNN baseline methods is demonstrated in Fig. 4. The primary variables are summarized in Table 1.
Figure 2: The left images are from a natural dehazing dataset, O-HAZE (Ancuti et al., 2018); the middle group is from a real-world surgical desmoking benchmark VASST-desmoke (Xia et al., 2025); and the line chart (right) shows the difference in error distributions between surgical desmoking and natural dehazing, computed from the entire dataset. The $\text{1}^{\text{st}}$ row displays the hazy or smoky images; the $\text{2}^{\text{nd}}$ row shows the smoke or haze mask estimated from the ground truth; and the $\text{3}^{\text{rd}}$ row presents the error magnitude map of the DCP prediction.
@@ 3.1.1 Physics-Based Modeling
The atmospheric scattering model (Narasimhan and Nayar, 2002, 2003; Fattal, 2008; Tan, 2008) is typically formulated as:
@@
where $c$ denotes a color channel; $x$ represents the spatial location of a pixel; $I\in\mathbb{R}^{H\times W\times C}$ is the observed image, where $C$, $W$, and $H$ are the number of color channels, width and height, respectively; $J$ indicates the corresponding haze-free image; $A\in\mathbb{R}^{1\times 1\times C}$ denotes the global atmospheric light; and $t\in\mathbb{R}^{H\times W\times 1}$ is the medium transmission. In particular, $t$ and $J$ can be further formulated as:
@@
where $\beta(\lambda)\in\mathbb{R}$ is the scattering coefficient for a specific wavelength $\lambda$ of light; $d\in\mathbb{R}^{H\times W\times 1}$ represents the scene depth; and $\rho\in\mathbb{R}^{H\times W\times C}$ is the normalized radiance of a scene point, which is mainly determined by reflectance of object surfaces (Narasimhan and Nayar, 2003; Tan, 2008).
Kaiming et al. (2011) proposed the Dark Channel Prior (DCP), which estimates the transmission map $t$ using the dark channel:
$\ \begin{aligned} t(x)&=1-D(x)\\ &=1-\underset{c\in C}{\mathrm{min}}\left(\underset{u\in\Omega(x;z)}{\mathrm{min}}\left(\frac{I(u,c)}{A(c)}\right)\right),\end{aligned}$ (4)
@@ 3.1.2 Deep Learning-Based Modeling
Clearly, in pixels where the smoke density is high and white or gray surgical instruments are present (i.e., $D\rightarrow 1$), the conventional physics-based model Eq. (5) may fail. Moreover, radial light attenuation can violate the assumption that $A\in\mathbb{R}^{1\times 1\times C}$; under such conditions, $A$ is more appropriately represented as $A\in\mathbb{R}^{H\times W\times C}$. Adhering to the original assumption may introduce more errors, whereas adopting the latter formulation significantly increases the degrees of freedom of the solution space. Therefore, Deep Neural Network (DNN), a powerful technique for non-linear mapping, is widely employed for smoke removal. Its formulation and training objective are typically expressed as:
@@
where $\Phi$ indicates a non-linear mapping parameterized by trainable weights $\theta$; $\theta^{*}$ represents the optimized weights obtained through training on the dataset $\mathcal{X}$; $\tilde{Z}$ and $\mathcal{L}$ denote priors and loss function, respectively, as determined by the specific method design. Nevertheless, its generalizability may be limited when deployed in unseen surgical scenarios.
@@ 3.2 Mixture-of-Experts and Optimization
Inspired by the Mixture-of-Experts (MoE), an effective mechanism for complementary advantages (Jacobs et al., 1991; Ding et al., 2025), we begin with an ideal gating model:
$\ \begin{aligned} J_{\mathrm{MoE}}(x,c)&=\tilde{W}(x)\cdot J_{\mathrm{DCP}}(x,c)\\ &\quad+(1-\tilde{W}(x))\cdot J_{\mathrm{DNN}}(x,c),\end{aligned}$ (8)
@@
$\ \begin{aligned} \varepsilon_{\mathrm{MoE}}(x,c)=W(x)\cdot\varepsilon_{\mathrm{DCP}}(x,c)\\ \quad+(1-W(x))\cdot\varepsilon_{\mathrm{DNN}}(x,c).\end{aligned}$ (12)
@@
$\ \begin{aligned} \min_{W(x)}\;\;&\mathbb{E}[\varepsilon_{\mathrm{MoE}}^{2}],\quad s.t.\;\;0\leq W(x)\leq 1.\end{aligned}$ (13)
Table 2: Comparison of Prediction Error Distributions with Gaussian and Laplacian Distributions via Jensen-Shannon (JS) Divergence in VASST-desmoke with real-world smoke and Cholec80 with synthetic smoke.
@@
DCP (Kaiming et al., 2011) 0.044989 0.013664 0.002782 0.010886
AODNet (Li et al., 2017) 0.077348 0.056151 0.006405 0.012263
CGAN-DC (Salazar-Colores et al., 2020) 0.033480 0.033894 0.007939 0.005278
De-smokeGCN (Chen et al., 2020) 0.027138 0.001988 0.007484 0.005802
GCANet (Chen et al., 2019) 0.009448 0.003621 0.003164 0.008474
LGUTransformer (Wang et al., 2024b) 0.010483 0.003129 0.016931 0.003623
MARS-GAN (Hong et al., 2023) 0.024441 0.064570 0.009878 0.005369
MSBDN (Dong et al., 2020) 0.020062 0.007690 0.005995 0.005536
RSTN (Wang et al., 2023) 0.026959 0.006900 0.012523 0.004193
SSIM-PAN (Sidorov et al., 2020) 0.018315 0.011999 0.005900 0.006517
SVPNet (Wang et al., 2024a) 0.022240 0.000675 0.007876 0.006492
@@
Figure 3: The Laplacian parameters (i.e., $\mu$ and $b$) estimated from the specific dataset, and indicator $W^{*}$ calculated with Eq. (19), are presented across different methods with the Confidence Interval (CI). The upper and lower rows correspond to Cholec80 and VASTT-desmoke, respectively.
@@
where $\mathcal{M}$ represents a method, either DCP or DNN. Collecting Eqs.(12), and (14), we have:
$\ \begin{aligned} \mathrm{Var}[\varepsilon_{\mathrm{MoE}}]&=2\cdot\big(W^{2}(x)\cdot b^{2}_{\mathrm{DCP}}(D(x))\\ &\quad+(1-W(x))^{2}\cdot b^{2}_{\mathrm{DNN}}(D(x))\big),\end{aligned}$ (15)
$\ \begin{aligned} \mathbb{E}[\varepsilon_{\mathrm{MoE}}]&=W(x)\cdot\mu_{\mathrm{DCP}}(D(x))\\ &\quad+(1-W(x))\cdot\mu_{\mathrm{DNN}}(D(x)).\end{aligned}$ (16)
Then, the original optimization problem Eq. (13) is rewritten as:
@@
We solve Eq. (17) and obtain the unconstrained solution:
@@
where $W^{*}(x;D)$ is the optimized $W(x)$ conditioned on the dark channel $D$ without constraints. For simplicity, $b_{\mathrm{DCP}}(D(x))$, $b_{\mathrm{DNN}}(D(x))$, $\mu_{\mathrm{DCP}}(D(x))$ and $\mu_{\mathrm{DNN}}(D(x))$ are written as $b_{1,x}$, $b_{2,x}$, $\mu_{1,x}$ and $\mu_{2,x}$, respectively. Subsequently, $W^{*}(x)$ is clipped to the interval $[0,1]$ to satisfy the constraint Eq. (18).
@@ 3.3 A New Restoration Formula
Following Eq. (19), Fig. 3 presents the trends of $W^{*}$ given $D$, based on statistics across all methods in two datasets: Cholec80 (Twinanda et al., 2017) with synthetic smoke, and the benchmark VASST-desmoke (Xia et al., 2024, 2025) containing real-world smoke. Deep learning models display varied trends, while a consistent negative correlation between $W^{*}$ and $D$ is observed. Therefore, in this work, $W$ is estimated by:
@@
Table 3: Pearson Correlation Coefficient between $W$ and $D$.
Method VASST-desmoke Cholec80
AODNet (Li et al., 2017) -0.065 -0.962
CGAN-DC (Salazar-Colores et al., 2020) -0.820 -0.963
De-smokeGCN (Chen et al., 2020) -0.763 -0.973
GCANet (Chen et al., 2019) -0.933 -0.765
LGUTransformer (Wang et al., 2024b) -0.899 -0.954
MARS-GAN (Hong et al., 2023) -0.977 -0.942
MSBDN (Dong et al., 2020) -0.899 -0.964
RSTN (Wang et al., 2023) -0.896 -0.945
SSIM-PAN (Sidorov et al., 2020) -0.842 -0.982
SVPNet (Wang et al., 2024a) -0.873 -0.965
Avg. -0.797 -0.942
@@
    We calculate Pearson Correlation Coefficient (PCC) between $W$ and $D$ for all methods. Most PCCs are close to 0.9-0.9, that strong negative linear correlation is obsereved, as reported in Table 3. For the low PCC of AODNet in VASST-desmoke, we observe that the origin AODNet struggles to deal with dense smoke (i.e., $D$ $\rightarrow$ 1), showing similar quantitative performance to DCP (i.e., $W$ $\rightarrow$ 0.5 means equal significance for DCP and AODNet); this is because AODNet is featured with only five convolutional layers and exhibits weaker reconstruction ability facing to dense smoke and fewer training data compared to Cholec80.
  2. 2.
@@
    Equation Eq. (20) is intuitive: a high intensity of dark channel (i.e., $D(x)\rightarrow 1$) statistically indicates dense smoke (Kaiming et al., 2011; Xia et al., 2025), which challenges DCP due to information loss and division by zero or very small values, thereby making DNN the dominant model (i.e., $W(x)\rightarrow 0$);
  4. 4.
@@
  5. 5.
    The denominator 1D1-D and the unknown parameter $A$ in Eq. (5) are eliminated in the subsequent derivation, stabilizing the computation, simplifying the modeling, and precluding hard value clipping.
By combining Eqs.(5), (8), and (20), we obtain:
$\ \begin{aligned} J_{\mathrm{MoE}}(x,c)&=I(x,c)-A(c)\cdot D(x)\\ &\quad+D(x)\cdot J_{\mathrm{DNN}}(x,c).\end{aligned}$ (21)
Then, we use the physical reflection model Eq. (3) to further decompose $J_{\mathrm{DNN}}$:
$\ \begin{aligned} J_{\mathrm{MoE}}(x,c)&=I(x,c)-A(c)\cdot D(x)\\ &\quad+D(x)\cdot A(c)\cdot\rho_{\mathrm{DNN}}(x,c),\end{aligned}$ (22)
where $\rho_{\mathrm{DNN}}(x,c)=J_{\mathrm{DNN}}(x,c)/A(c)$ represents the normalized radiance obtained from Eq. (3). Equation Eq. (22) can be reformulated as:
$\ \begin{aligned} J_{\mathrm{MoE}}(x,c)&=I(x,c)-D(x)\cdot A(c)\cdot(1-\rho_{\mathrm{DNN}}(x,c))\\ &=I(x,c)-\mathcal{D}(x)\cdot(1-\rho_{\mathrm{DNN}}(x,c)),\end{aligned}$ (23)
where $\mathcal{D}$ denotes the denormalized dark channel $D$. Recalling Eq. (4), $\mathcal{D}$ is defined as:
@@
where $A$ is eliminated, thereby avoiding the challenging estimation of global light illuminance in laparoscopic surgical scenes. In this case, the original DNN modeling in Eqs.(6) and (7) can be improved as:
@@
where $\sigma$ is a differentiable function (typically a sigmoid) that maps the DNN output to the interval $[0,1]$. It can be ignored if the output is already constrained within this range, depending on the specific model design. The proposed formulation introduces only minimal modifications to the original DNN model $\Phi_{\theta}$, without adding any additional trainable parameters. Notably, SurgiATM leverages the denormalized dark channel as a key element, in line with recent methods (Chen et al., 2019; Hong et al., 2023; Chen et al., 2020), suggesting that the dark channel prior plays a useful role in surgical desmoking (Xia et al., 2025).
@@ 3.4 Gradient Computation and Refinement
In this section, we compute the gradient with respect to the model prediction $\rho_{\mathrm{DNN}}$ for two commonly used loss functions: Mean Absolute Error (MAE or L1) loss and Mean Squared Error (MSE or L2) loss. Their gradients are denoted as $\nabla_{\rho}\mathcal{L}_{1}$ and $\nabla_{\rho}\mathcal{L}_{2}$, respectively:
$\ \nabla_{\rho}\mathcal{L}_{1}=\begin{cases}-\mathcal{D},&J>1-\mathcal{D}\cdot(1-\rho);\\ \mathcal{D},&J<1-\mathcal{D}\cdot(1-\rho);\\ 0,&J=1-\mathcal{D}\cdot(1-\rho),\end{cases}$ (27)
$\ \nabla_{\rho}\mathcal{L}_{2}=-\mathcal{D}\cdot(J-(1-\mathcal{D}\cdot(1-\rho))),$ (28)
@@
where $\eta$ is a smoothing factor; $\eta=0$ corresponds to no refinement; and $\eta\rightarrow+\infty$ implies $\hat{\mathcal{D}}\rightarrow 1$, causing Eq. (26) to degenerate into conventional residual prediction. Ablation results for $\eta$ and $z$ are reported in Section 5.1.
Figure 4: (a): General baseline framework for DNN-based desmoking methods, and the dashed box represents the part of discriminator for GAN-based methods. (b): General framework for using SurgiATM. The modification does not involve main architecture (i.e., DNN models), suggesting minimal adjustment overhead.
3.5 Plug into Baselines
SurgiATM can be regarded as a processing pipeline (or output layer) appended to the end of DNN models (see Fig. 4, Fig. 4). It introduces only minimal modification to the original baseline. In conventional settings, the output of a DNN model is directly treated as desmoked results, based on which losses are computed and gradients are propagated during training (see Fig. 4(a)).
With SurgiATM, the DNN output is first interpreted as a normalized radiance $\rho_{\text{DNN}}$, and then passed into SurgiATM. Losses are computed using the reconstructed results produced by SurgiATM, and gradients are backpropagated through SurgiATM to the baseline DNN models. For GAN-based approaches, after incorporating SurgiATM, the discriminator should no longer operate directly on the final output of generator; instead, it should take the SurgiATM-processed results as input (see Fig. 4(b)). In particular, for GAN-based baselines that pass intermediate feature maps from the generator to the discriminator, SurgiATM should not be applied to these intermediate representations or used as a shortcut, as it is specifically designed to accept normalized radiance with physical meaning.
On the other hand, Diffusion-based methods have recently emerged as a prevalent paradigm in generative AI, and have also been explored for the surgical desmoking task. However, SurgiATM does not currently support diffusion-based approaches. This is because SurgiATM is developed for methods that directly predict smokeless images in a single forward pass and is derived from classic physical models that formulate a direct mapping from smoky frames to smokeless ones. This formulation is fundamentally different from Diffusion models, in which DNNs are charged to predicting Gaussian noise components to be removed (i.e., learning a mapping from noisy images to Gaussian noise) through an iterative denoising process involving multiple inference steps. In addition, comparatively higher computational cost of Diffusion models remains a challenge for real-time deployment (Kazerouni et al., 2023). Consequently, Diffusion-based methods are beyond the scope of this work.
Table 4: Training and evaluation on VASST-desmoke (real smoke). There is no overlap videos between training and evaluation sets. The results are represented in the form of ”Baseline / Baseline + SurgiATM”.
Method Backbone Loss Function CIEDE2000 $\downarrow$ PSNR $\uparrow$ RMSE $\downarrow$ SSIM $\uparrow$
AODNet (Li et al., 2017) CNN MSE 10.349 / 6.327 17.378 / 21.994 0.147 / 0.090 0.693 / 0.789
GCANet (Chen et al., 2019) CNN MSE 7.893 / 5.369 21.091 / 23.907 0.095 / 0.073 0.729 / 0.806
De-smokeGCN (Chen et al., 2020) U-Net MAE + Others 9.577 / 7.348 19.779 / 22.127 0.115 / 0.094 0.550 / 0.717
MSBDN (Dong et al., 2020) U-Net MSE 6.679 / 5.164 21.604 / 23.881 0.088 / 0.073 0.651 / 0.813
SVPNet (Wang et al., 2024a) U-Net MSE + Others 5.574 / 5.141 23.464 / 24.362 0.073 / 0.070 0.789 / 0.820
CGAN-DC (Chen et al., 2019) cGAN Adversarial + Others 6.735 / 5.812 22.453 / 23.809 0.083 / 0.079 0.676 / 0.782
MARS-GAN (Hong et al., 2023) CycleGAN Adversarial + Others 6.389 / 5.899 20.121 / 23.488 0.113 / 0.076 0.602 / 0.819
SSIM-PAN (Sidorov et al., 2020) GAN Adversarial + Others 7.046 / 5.811 21.465 / 23.477 0.090 / 0.075 0.655 / 0.750
RSTN (Wang et al., 2023) SwinT Charbonnier + Others 5.385 / 4.976 23.263 / 24.306 0.074 / 0.071 0.750 / 0.824
LGUTransformer (Wang et al., 2024b) SwinT + U-Net Charbonnier 9.927 / 5.990 18.238 / 23.376 0.121 / 0.080 0.510 / 0.761